Determine whether u and v are orthogonal vectors,make an acute or obtuse angle:

1.u=<1,3,-2>; v=<-5,3,2>

2.u=<1,-2,4>; v=<5,3,7>


1
Expert's answer
2021-06-18T08:40:33-0400

1.u=<1,3,-2>; v=<-5,3,2>



The vectors are


u = i + 3j - 2k

v = -5i +3j + 2k


"\cos \theta" = "\dfrac{u.v}{|u||v|}"


cosθ = "\dfrac{(i + 3j - 2k ) . (-5i +3j + 2k) }{\sqrt{14}*\sqrt{38}}"


cosθ = -"\dfrac{- 5 + 9 - 4}{\sqrt{14}*\sqrt{38}}" = 0


"\theta" = "\dfrac{\pi}{2}"



Hence the vectors u and v are orthogonal to each other.




2.u=<1,-2,4>; v=<5,3,7>


The vectors are


u = i - 2j + 4k

v = 5i +3j + 7k


"\cos \theta" = "\dfrac{u.v}{|u||v|}"


cosθ = "\dfrac{(i - 2j + 4k ) . (5i +3j + 7k) }{\sqrt{21}*\sqrt{83}}"


cosθ = "\dfrac{5 - 6 + 28}{\sqrt{21}*\sqrt{83}}"


"\cos \theta" = "\dfrac{27}{\sqrt{21}*\sqrt{83}}"


"\cos\theta" = 0.647


"\theta" = "49.68\degree"



Since, "\theta" < 90"\degree" so u and v make the acute angle with each other.